OpenStudy (anonymous):

-tan^2 x + sec^2 x = 1

OpenStudy (cwrw238):

to prove this you just need to rearrange the standard identity sec^2x = 1 + tan^2 x subtract tan^2 x from both sides: -tan^2 x + sec^2 x = 1 + tan^2 x - tan^2 x -tan^2 x + sec^2 x = 1

OpenStudy (anonymous):

OK :) got it

OpenStudy (cwrw238):

if you want another way to do this: -tan^2 x + sec^2 x = - sin^2 x + 1 ------- ------ cos^2 x cos^2 x = 1 - sin^2 x ----------- cos^2 x = cos^2 x ------- = 1 (because sin^2 x + cos^2 x = 1) cos^2 x

OpenStudy (anonymous):

Wow thank you, yea I like this one better :)